# EP240414a: A Gamma-Ray Burst Jet Weakened by an Extended Circumstellar Material

### v1

@@ Abstract
The recent Einstein Probe (EP) event EP240414a exhibits several unusual observational features. Its prompt and afterglow emissions place it between long gamma-ray bursts (LGRBs) and low-luminosity GRBs (LLGRBs). The event is followed by a fast optical transient (AT 2024gsa), initially exhibiting a thermal-like spectrum but later evolving into an unusually red peak at $\sim 3-5$ days, which is difficult to explain with thermal emission. Using our generalized analytic framework for jet propagation in a circumstellar material (CSM; Hamidani et al. 2025), we explore a scenario in which a conventional LGRB jet is launched in a progenitor surrounded by a dense CSM. For a CSM of $\sim 0.03M_{\odot}$ extending to $\sim 3\times 10^{13}$ cm, we find that the jet is significantly weakened before breaking out, becoming “barely failed”, an intermediate state between successful (LGRB) and completely failed (LLGRB) jets. This scenario naturally explains EP240414a’s multi-wavelength observations, with the early thermal component produced by cocoon cooling emission, and the red peak explained by non-thermal afterglow emission from the mildly relativistic barely failed jet (and its inner cocoon). Our work demonstrates the important role of extended CSM in shaping GRB jets and illustrates how early multi-wavelength follow-up observations can reveal the physically diverse nature of jet-driven transients.
Gamma-ray bursts (629), Relativistic jets (1390), Hydrodynamics (1963)
@@ 1 Introduction
In contrast, low-luminosity GRBs (LLGRBs) exhibit lower isotropic energies ($E_{\rm iso}\sim 10^{48}-10^{50}$ erg) and softer spectra (e.g., GRB 060218 and GRB 100316D; Campana et al. 2006; Starling et al. 2011). LLGRBs are thought to be explained by shock breakout emission, implying that they do not require relativistic jets (Waxman et al. 2007; Nakar & Sari 2012). In particular, it has been suggested that the weakness of prompt emission in LLGRBs (relative to LGRBs) can be explained by a dense circumstellar medium (CSM) around the progenitor, which prevents the jet from breaking out and produces a non-relativistic cocoon that powers shock breakout emission as it breaks out of the CSM (Nakar 2015).
@@
Studies of the interaction between LGRB jets and CSM suggest that the CSM can not only choke the jet, but also produce bright cooling emission from shocked CSM (i.e., cocoon), leading to transients in the form of fast blue optical transients (FBOTs; Nakar 2015; Margutti et al. 2019; Suzuki & Maeda 2022; Suzuki et al. 2024; also see Gottlieb et al. 2022).
@@
The recent Einstein Probe (EP) mission, with its sensitivity to soft X-ray, has identified numerous X-ray transients, including EP240414a (Sun et al. 2024). EP240414a exhibits an intermediate intensity between LLGRBs and LGRBs in both prompt and afterglow emissions (Sun et al. 2024; van Dalen et al. 2024; Srivastav et al. 2025; Bright et al. 2024). Here, assuming that certain WR stars possess a CSM and that EP240414a originates from a conventional collapsar event, we explore the necessary CSM conditions to explain EP240414a’s unusual features. We show that with a CSM mass of $\sim 0.03M_{\odot}$ extending to $\sim 3\times 10^{13}$ cm, the LGRB jet becomes barely failed, providing a scenario that coherently explains EP240414a’s unique characteristics.
@@ 2 Observational properties
EP240414a is an X-ray transient detected by EP (Sun et al. 2024) and subsequently observed across the X-ray, radio, optical and infrared (IR) bands. In the optical and IR bands, a luminous “FBOT” counterpart (AT 2024gsa, although red in color) was discovered, which evolved into a BL SN Ic at later times (SN 2024gsa, van Dalen et al. 2024; Srivastav et al. 2025; Bright et al. 2024). The event was located at $z=0.401$ ($D_{L}=2245$ Mpc), at a large offset of $26$ kpc from the spiral galaxy J1246 (van Dalen et al. 2024; Sun et al. 2024). It is unclear whether it is physically offset or hosted by an unresolved faint-satellite galaxy.
@@
    Phase I ($\lesssim 1$ day): Early follow-up observations found a luminous FBOT (AT 2024gsa) consisting of two parts (phase I and phase II). Spectroscopy in optical/IR during phase I (at $\sim 0.6$ day) shows a clear blue spectral slope ($F_{\nu}\propto\nu^{0.9}$) that is inconsistent with typical afterglow emission ($F_{\nu}\propto\nu^{-1}$; see Section 3.3 in van Dalen et al. 2024).
@@
    Phase II ($\sim 1-10$ days): AT 2024gsa peaked with an absolute magnitude of $\sim-21$ mag (in r-band, at $\sim 4$ days). At its peak, it exhibited a red color that challenges thermal emission scenarios (Sun et al. 2024; van Dalen et al. 2024). Srivastav et al. (2025) pointed out that a thermal model would require a photospheric velocity of $\sim 0.7c$ to reproduce the color; which in turn would introduce additional issues due to relativistic effects on observer time and color shift. Spectroscopy in optical/IR around the peak (at $\sim 3-5$ days) shows a flat featureless spectrum, which is more compatible with afterglow scenarios (see Figure 2 in van Dalen et al. 2024).
@@
    Phase III ($\gtrsim 10$ days): A BL Type Ic supernova, SN 2024gsa, was spectroscopically identified, exhibiting similarities to other LGRB-SNe such as SN 1998bw and SN 2006aj (Sun et al. 2024; van Dalen et al. 2024). However, SN 2024gsa appears significantly redder than typical BL Ic SNe (Taddia et al. 2019; see Appendix A).
@@
    Radio emission (days–months): The emission is consistent with the lower luminosity end of LGRB afterglows, yet it remains orders of magnitude brighter than that of LLGRBs (Sun et al. 2024; Bright et al. 2024). This rules out the scenario of a completely failed jet and disfavors the presence of an energetic, highly relativistic jet (Bright et al. 2024).
First, the observation in 4) points to an ordinary collapsar event involving a massive star (Woosley 1993; MacFadyen & Woosley 1999). From the properties in 1) and 5), we deduce that a completely choked/failed jet is ruled out. Additionally, from 1) and 5), we deduce that a standard energetic LGRB jet is also disfavored. From 2) and 3), we deduce that: although phase I of AT 2024gsa is thermal and could possibly be powered by the jet/SN CSM interaction (see van Dalen et al. 2024); phase II of AT 2024gsa is inconsistent with thermal emission and could be afterglow powered (Srivastav et al. 2025), which would be consistent with 5).
Based on these observations, we propose an intermediate scenario, where the jet is neither completely failed (as in LLGRBs) nor successful (as in LGRBs). We propose a “barely failed” jet scenario to explain EP240414a/AT 2024gsa, where the jet is significantly weakened (energetically and relativistically) before breaking out (see Eisenberg et al. 2022 and their Figure 6; also see Hamidani et al. 2025). We consider that this weakening is due to an extended CSM around the progenitor (similar but not identical to CSM proposed for LLGRBs in Nakar 2015; see Hamidani et al., 2025 for details on parameter space for this scenario) which also conveniently explains the early thermal emission ($\lesssim 1$ day; van Dalen et al. 2024).
@@ 3.1 The Model
We follow the same approximations as in Hamidani et al. (2025) with respect to the CSM and the jet. In particular, we consider a conventional LGRB jet. We consider a total energy budget of $E_{\rm{eng}}\sim 10^{52}$ erg deposited on a timescale $t_{\rm{eng}}=100$ s (which is comparable to the free-fall timescale of the WR progenitor star; Matzner 2003) in the form of a relativistic jet with an initial opening angle $\theta_{0}=10^{\circ}$. For the CSM we assume a radius $R_{\rm CSM}=3\times 10^{13}$ cm and a mass $M_{\rm CSM}=0.03M_{\odot}$ so that the jet is barely failed (see Hamidani et al. 2025) and thermal emission is consistent with observations (van Dalen et al. 2024).
@@ 3.4.3 The SN
For the SN component (SN 2024gsa), based on its similarity to SN 1998bw (Sun et al. 2024; van Dalen et al. 2024), and for simplicity, we use the template of SN 1998bw in Clocchiatti et al. (2011) at $z=0.401$. This is a crude assumption, as a detailed study of the SN is beyond the scope of this work.
From the analysis of SN 2024gsa and its comparison with other BL Ic SNe (Taddia et al. 2019; also Prentice et al. 2016; Stritzinger et al. 2018), we found evidence of moderate host extinction, which is also consistent with the high column density measured in X-ray. Our analysis indicates $E(B-V)\sim 0.15-0.3$. In the following, we adopt a conservative value $E(B-V)\sim 0.15$ (see Appendix A).
Figure 2: Multi-band observations of EP240414a’s counterparts and the theoretical interpretation. Observed fluxes in X-ray, g-band, r-band, i-band, z-band, and radio are shown (black, purple, blue, green, orange, and red, respectively; circles for detections and downward triangles for upper limits) in the observer-frame (bottom) and in the rest-frame (top) time (for $z=0.401$). Estimated total flux contributions (solid lines) along with individual components are shown: afterglow emission from the mildly relativistic jet (dashed) and the inner-cocoon (dashed dotted) [see Section 3.4.2]; cooling emission from the non-relativistic outer-cocoon (dotted) [see Section 3.4.1]; and the SN component (spaced-dotted) [see Section 3.4.3]. The three phases are indicated with the corresponding dominant contributions (as in Sun et al. 2024). Photometric Data is taken from Sun et al. (2024); van Dalen et al. (2024); Srivastav et al. (2025) and has been corrected for Galactic extinction [$E(B-V)=0.032$]. In addition, moderate host extinction [$E(B-V)=0.15$] is assumed based on the analysis of the SN color and the X-ray column density (see Appendix A). The X-ray data are taken from Table 4 in Sun et al. (2024). Radio data are taken from Table 1 in Bright et al. (2024).
@@ 4.1 Cooling emission (phase I)
In Figure 2, our model of cooling emission (dotted line; phase I) is consistent with optical photometric data (both flux and color) at $\sim 0.1-1$ day. Our model predicts high-temperature thermal emission during this phase, which aligns with the very blue optical spectrum observed at $\sim 0.6$ days (see Section 3.3 in van Dalen et al. 2024), coinciding with the decay phase of this cooling emission. At that time, the thermal component dominates at shorter wavelengths (e.g., g-band) due to its high temperature ($\sim 3\times 10^{4}$ K), while at longer wavelengths (e.g., i-band), it contributes comparably to the afterglow. This naturally explains the observed blue spectral slope and its evolution at later times (see Figure 2 in van Dalen et al. 2024). Thus, our scenario provides a consistent interpretation of both the photometric and spectroscopic data in phase I.
@@ 4.2 Afterglow emission (phases I – III)
General parameters are: $\theta_{v}=0$ rad, $n_{0}=0.6$ cm-3, and $p=2.4$.
Our model shows that, in the optical/IR bands, afterglow emission transitions from being subdominant in phase I to becoming the dominant component in phase II (the FBOT peak; AT 2024gsa). This is consistent with spectroscopic observations of phase I, which suggest a thermal origin (van Dalen et al. 2024), as our afterglow component (non-thermal) remains subdominant. Furthermore, this aligns with the findings of Srivastav et al. (2025), which indicate that the peak of phase II is inconsistent with a thermal origin, as it would require a very large photosphere (moving with $\sim 0.7c$) to match the color. Overall, our model, despite its simplicity, provides a reasonable explanation for the photometric data across all bands, with the exception of the transition between phase I and phase II. We suspect that this is due to our reliance on only two jet components [^2] In reality, the jet is expected to exhibit a multi-component structure (e.g., Gottlieb et al. 2021), and our model lacks an intermediate component that would peak between the jet and the inner-cocoon, which would improve the fit in this transition phase..
@@ 4.3 The Supernova (phase III)
In Figure 2, we show the template of SN 1998bw from Clocchiatti et al. (2011) as a reference (spaced-dotted line; phase III; assumed to be powered by Ni56 decay). Apart from the g-band, where the observed data indicate a significantly dimmer SN, this template provides a reasonable fit. The discrepancy in the g-band is likely due to a combination of two factors; to a lesser extent, uncertainties in the intrinsic luminosity of SN 2024gsa, which appears intermediate between SN 1998bw and SN 2006aj (Sun et al. 2024; van Dalen et al. 2024); and to a greater extent, uncertainties in the extinction at the rest-frame of EP240414a (our assumption of $E(B-V)\sim 0.15$ is conservative; see Appendix A). Although the nature of SN 2024gsa is intriguing, it is beyond the scope of this study, which focuses on jet-driven transients evolving on timescales of $<10$ days.

### v3

@@ Abstract
The recent Einstein Probe (EP) event EP240414a exhibits several unusual observational features. Its prompt and afterglow emissions place it between long gamma-ray bursts (LGRBs) and low-luminosity GRBs (LLGRBs). The event is followed by a fast optical transient (AT 2024gsa), initially exhibiting a thermal-like spectrum but later evolving into an unusually red peak at $\sim 3-5$ days, which is difficult to explain with thermal emission. Using our generalized analytic framework for jet propagation in a circumstellar material (CSM; Hamidani et al. 2025), we explore a scenario in which a conventional LGRB jet is launched in a progenitor surrounded by a dense CSM. For a CSM of $\sim 0.03M_{\odot}$ extending to $\sim 3\times 10^{13}$ cm, we find that the jet is significantly weakened before breaking out, becoming “barely failed”, an intermediate state between successful (LGRB) and completely failed (LLGRB) jets. This scenario naturally explains EP240414a’s multi-wavelength observations, with the early thermal component produced by cocoon cooling emission, and the red peak explained by non-thermal afterglow emission from the mildly relativistic barely failed jet (and its inner-cocoon). Our work demonstrates the important role of extended CSM in shaping GRB jets and illustrates how early multi-wavelength follow-up observations can reveal the physically diverse nature of jet-driven transients.
Gamma-ray bursts (629), Relativistic jets (1390), Circumstellar matter(241), Hydrodynamics (1963), Transient sources (1851), X-ray transient sources (1852), Core-collapse supernovae(304), Type Ic supernovae(1730), High-energy astrophysics (739)
@@ 1 Introduction
In contrast, low-luminosity GRBs (LLGRBs) exhibit lower isotropic energies ($E_{\rm iso}\sim 10^{48}-10^{50}$ erg) and softer spectra (e.g., GRB 060218 and GRB 100316D; Campana et al. 2006; Soderberg et al. 2006; Starling et al. 2011). LLGRBs are thought to be explained by shock breakout emission, implying that they do not require relativistic jets (Waxman et al. 2007; Nakar & Sari 2012). In particular, it has been suggested that the weakness of prompt emission in LLGRBs (relative to LGRBs) can be explained by a dense circumstellar medium (CSM) around the progenitor, which prevents the jet from breaking out and produces a non-relativistic cocoon that powers shock breakout emission as it breaks out of the CSM (Nakar 2015; Irwin & Chevalier 2016; Irwin & Hotokezaka 2024).
@@
Studies of the interaction between LGRB jets and CSM suggest that the CSM can not only choke the jet, but also produce bright cooling emission from shocked CSM (i.e., cocoon), leading to transients in the form of fast blue optical transients (FBOTs; Nakar 2015; Margutti et al. 2019; Suzuki & Maeda 2022; Suzuki et al. 2024; also see Gottlieb et al. 2022; and Duffell & Ho 2020).
@@
The recent Einstein Probe (EP) mission, with its sensitivity to soft X-ray, has identified numerous X-ray transients, including EP240414a (Sun et al. 2024). EP240414a exhibits an intermediate intensity between LLGRBs and LGRBs in both prompt and afterglow emissions (Sun et al. 2024; van Dalen et al. 2025; Srivastav et al. 2025; Bright et al. 2024). Here, assuming that certain WR stars possess a CSM and that EP240414a originates from a conventional collapsar event, we explore the necessary CSM conditions to explain EP240414a’s unusual features. We show that with a CSM mass of $\sim 0.03M_{\odot}$ extending to $\sim 3\times 10^{13}$ cm, the LGRB jet becomes barely failed, providing a scenario that coherently explains EP240414a’s unique characteristics.
@@ 2 Observational properties
EP240414a is an X-ray transient detected by EP (Sun et al. 2024) and subsequently observed across the X-ray, radio, optical and infrared (IR) bands. In the optical and IR bands, a luminous “FBOT” counterpart (AT 2024gsa, although red in color) was discovered, which evolved into a BL SN Ic at later times (SN 2024gsa, van Dalen et al. 2025; Srivastav et al. 2025; Bright et al. 2024). The event was located at $z=0.401$ ($D_{\rm L}\approx 2245$ Mpc; assuming a $\Lambda$CDM cosmology with a Hubble constant $H_{0}=67.7\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{\mathrm{M}}=0.309$, and $\Omega_{\Lambda}=0.691$; Planck Collaboration et al. 2016), at a large offset of $26$ kpc from the spiral galaxy J1246 (van Dalen et al. 2025; Sun et al. 2024). It is unclear whether it is physically offset or hosted by an unresolved faint-satellite galaxy.
@@
    Phase I ($\lesssim 1$ day): Early follow-up observations found a luminous FBOT (AT 2024gsa) consisting of two parts (phase I and phase II). Spectroscopy in optical/IR during phase I (at $\sim 0.6$ day) shows a clear blue spectral slope ($F_{\nu}\propto\nu^{0.9}$) that is inconsistent with typical afterglow emission ($F_{\nu}\propto\nu^{-1}$; see Section 3.3 in van Dalen et al. 2025).
@@
    Phase II ($\sim 1-10$ days): AT 2024gsa peaked with an absolute magnitude of $\sim-21$ mag (in r-band, at $\sim 4$ days). At its peak, it exhibited a red color that challenges thermal emission scenarios (Sun et al. 2024; van Dalen et al. 2025). Srivastav et al. (2025) pointed out that a thermal model would require a photospheric velocity of $\sim 0.7c$ to reproduce the color; which in turn would introduce additional issues due to relativistic effects on observer time and color shift. Spectroscopy in optical/IR around the peak (at $\sim 3-5$ days) shows a flat featureless spectrum, which is more compatible with afterglow scenarios (see Figure 2 in van Dalen et al. 2025).
@@
    Phase III ($\gtrsim 10$ days): A BL Type Ic supernova, SN 2024gsa, was spectroscopically identified, exhibiting similarities to other LGRB-SNe such as SN 1998bw and SN 2006aj (Sun et al. 2024; van Dalen et al. 2025). However, SN 2024gsa appears significantly redder than typical BL Ic SNe (Taddia et al. 2019; see Appendix A).
@@
    Radio emission (days–months): The emission is consistent with the lower luminosity end of LGRB afterglows, viewed on-axis, yet it remains orders of magnitude brighter than that of LLGRBs (Sun et al. 2024; Bright et al. 2024). This rules out the scenario of a completely failed jet, and disfavors the presence of an energetic highly relativistic jet (Bright et al. 2024).
First, the observation in 4) points to an ordinary collapsar event involving a massive star (Woosley 1993; MacFadyen & Woosley 1999). From the properties in 1) and 5), we deduce that a completely choked/failed jet is ruled out. Additionally, from 1) and 5), we deduce that a standard energetic LGRB jet is also disfavored. From 2) and 3), we deduce that: although phase I of AT 2024gsa is thermal and could possibly be powered by the jet/SN CSM interaction (see van Dalen et al. 2025); phase II of AT 2024gsa is inconsistent with thermal emission and could be afterglow powered (Srivastav et al. 2025), which would be consistent with 5).
Based on these observations, we propose an intermediate scenario, where the jet is neither completely failed (as in LLGRBs) nor successful (as in LGRBs). We propose a “barely failed” jet scenario to explain EP240414a/AT 2024gsa, where the jet is significantly weakened (energetically and relativistically) before breaking out (see Eisenberg et al. 2022 and their Figure 6; also see Hamidani et al. 2025). We consider that this weakening is due to an extended CSM around the progenitor (similar but not identical to CSM proposed for LLGRBs in Nakar 2015; see Hamidani et al. 2025 for details on parameter space for this scenario) which also conveniently explains the early thermal emission ($\lesssim 1$ day; van Dalen et al. 2025).
@@ 3.1 The Model
We follow the same approximations as in Hamidani et al. (2025) with respect to the CSM and the jet. In particular, we consider a conventional LGRB jet. We consider a total energy budget of $E_{\rm{eng}}\sim 10^{52}$ erg deposited on a timescale $t_{\rm{eng}}=100$ s (which is comparable to the free-fall timescale of the WR progenitor star; Matzner 2003) in the form of a relativistic jet with an initial opening angle $\theta_{0}=10^{\circ}$. For the CSM we assume a radius $R_{\rm CSM}=3\times 10^{13}$ cm and a mass $M_{\rm CSM}=0.03M_{\odot}$ so that the jet is barely failed (see Hamidani et al. 2025) and thermal emission is consistent with observations (van Dalen et al. 2025).
@@ 3.4.3 The SN
For the SN component (SN 2024gsa), based on its similarity to SN 1998bw (Sun et al. 2024; van Dalen et al. 2025), and for simplicity, we use the template of SN 1998bw in Clocchiatti et al. (2011) at $z=0.401$. This is a crude assumption, as a detailed study of the SN is beyond the scope of this work.
From the analysis of SN 2024gsa and its comparison with other BL Ic SNe (Taddia et al. 2019; also Prentice et al. 2016; Stritzinger et al. 2018; Srinivasaragavan et al. 2024), we found evidence of moderate host extinction, which is also consistent with the high column density measured in X-ray. Our analysis indicates $E(B-V)\sim 0.15-0.3$. In the following, we adopt a conservative value $E(B-V)\sim 0.15$ (see Appendix A).
Figure 2: Multi-band observations of EP240414a’s counterparts and the theoretical interpretation. Observed fluxes in X-ray, g-band, r-band, i-band, z-band, and radio are shown (black, purple, blue, green, orange, and red, respectively; circles for detections and downward triangles for upper limits) in the observer-frame (bottom) and in the rest-frame (top) time (for $z=0.401$). Estimated total flux contributions (solid lines) along with individual components are shown: afterglow emission from the mildly relativistic jet (dashed) and the inner-cocoon (dashed dotted) [see Section 3.4.2]; cooling emission from the non-relativistic outer-cocoon (dotted) [see Section 3.4.1]; and the SN component (spaced-dotted) [see Section 3.4.3]. The three phases are indicated with the corresponding dominant contributions (as in Sun et al. 2024). Photometric Data is taken from Sun et al. (2024); van Dalen et al. (2025); Srivastav et al. (2025) and has been corrected for Galactic extinction [$E(B-V)=0.032$]. In addition, moderate host extinction [$E(B-V)=0.15$] is assumed based on the analysis of the SN color and the X-ray column density (see Appendix A). The X-ray data are taken from Table 4 in Sun et al. (2024). Radio data are taken from Table 1 in Bright et al. (2024). A summary of the model parameters is given in Table 1 in Appendix B.
@@ 4.1 Cooling emission (phase I)
In Figure 2, our model of cooling emission (dotted line; phase I) is consistent with optical photometric data (both flux and color) at $\sim 0.1-1$ day. Our model predicts high-temperature thermal emission during this phase, which aligns with the very blue optical spectrum observed at $\sim 0.6$ days (see Section 3.3 in van Dalen et al. 2025), coinciding with the decay phase of this cooling emission. At that time, the thermal component dominates at shorter wavelengths (e.g., g-band) due to its high temperature ($\sim 3\times 10^{4}$ K), while at longer wavelengths (e.g., i-band), it contributes comparably to the afterglow. This naturally explains the observed blue spectral slope and its evolution at later times (see Figure 2 in van Dalen et al. 2025). Thus, our scenario provides a consistent interpretation of both the photometric and spectroscopic data in phase I.
@@ 4.2 Afterglow emission (phases I – III)
General parameters are: $\theta_{v}=0$ rad, $n_{0}=0.6$ cm-3, and $p=2.4$ (also see Table 1 in Appendix B).
Our model shows that, in the optical/IR bands, afterglow emission transitions from being subdominant in phase I to becoming the dominant component in phase II (the FBOT peak; AT 2024gsa). This is consistent with spectroscopic observations of phase I, which suggest a thermal origin (van Dalen et al. 2025), as our afterglow component (non-thermal) remains subdominant. Furthermore, this aligns with the findings of Srivastav et al. (2025), which indicate that the peak of phase II is inconsistent with a thermal origin, as it would require a very large photosphere (moving with $\sim 0.7c$) to match the color. Overall, our model, despite its simplicity, provides a reasonable explanation for the photometric data across all bands, with the exception of the transition between phase I and phase II. We suspect that this is due to our reliance on only two jet components [^2] In reality, the jet is expected to exhibit a multi-component structure (e.g., Gottlieb et al. 2021), and our model lacks an intermediate component that would peak between the jet and the inner-cocoon, which would improve the fit in this transition phase.. Allowing for a variation in the viewing angle near the jet axis would improve the fit across the transition phase. In Zheng et al. (2025) a different viewing angle is considered, and phase II (its rise in particular) is explained by a top-hat relativistic jet viewed with $\theta_{v}\sim 10-15^{\circ}$.
@@ 4.3 The Supernova (phase III)
In Figure 2, we show the template of SN 1998bw from Clocchiatti et al. (2011) as a reference (spaced-dotted line; phase III; assumed to be powered by Ni56 decay). Apart from the g-band, where the observed data indicate a significantly dimmer SN, this template provides a reasonable fit. The discrepancy in the g-band is likely due to a combination of two factors; to a lesser extent, uncertainties in the intrinsic luminosity of SN 2024gsa, which appears intermediate between SN 1998bw and SN 2006aj (Sun et al. 2024; van Dalen et al. 2025); and to a greater extent, uncertainties in the extinction at the rest-frame of EP240414a (our assumption of $E(B-V)\sim 0.15$ is conservative; see Appendix A). Although the nature of SN 2024gsa is intriguing, it is beyond the scope of this study, which focuses on jet-driven transients evolving on timescales of $<10$ days.